OpenStudy (anonymous):

@Mathematics factor out the GCF from the polynomial 1/6x^7+5/6x^4-7/6x^6+1/6x

OpenStudy (anonymous):

looks like common factor is \[\frac{1}{6}x\]

OpenStudy (anonymous):

but how do you find it?

OpenStudy (anonymous):

\[\frac{1}{6}x^7+\frac{5}{6}x^4-\frac{7}{6}x^6+\frac{1}{6}x\] \[\frac{1}{6}x(x^6+5x^3-7x^5+1)\]

OpenStudy (anonymous):

well each term has a 6 in the denominator, and every term has an x in it, so you can "factor it out"

OpenStudy (anonymous):

oh okay i get it now thank you so much